2. A boy is flying a kite at a height of 150 ft. If the kite moves horizontally away from the


boy at the rate of 20 ft/sec, how fast is the string being paid out when the kite is 250 ft from him?

1
Expert's answer
2021-12-20T13:35:59-0500


Pythagorean Theorem


"L^2=x^2+150^2"

Differentiate both sides with respect to "t"


"2L(\dfrac{dL}{dt})=2x(\dfrac{dx}{dt})"

Solve for "\dfrac{dL}{dt}"


"\dfrac{dL}{dt}=(\dfrac{x}{L})\dfrac{dx}{dt}"

Substitute "L=\sqrt{x^2+150^2}"


"\dfrac{dL}{dt}=(\dfrac{x}{\sqrt{x^2+150^2}})\dfrac{dx}{dt}"

Given "\dfrac{dx}{dt}=20 ft/sec"

If "x=250\ ft"


"\dfrac{dL}{dt}=(\dfrac{250}{\sqrt{250^2+150^2}})(20)"

"\dfrac{dL}{dt}=\dfrac{100}{\sqrt{34}}\ ft/sec\approx17.15\ ft/s"